Prev Next

SW&IT Meeting Minutes 2020-08-31

Attendees: Carl, David, Igor, Matthieu, Trey, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200824
LR20200824 - Bec, Matthieu 40.0
LR20200824 - Holmberg, Carl 40.0
LR20200824 - Edmeades, David 44.0
LR20200824 - Sola, Igor 0.0
LR20200824 - Hooper, Stephen 0.0
LR20200824 - Bryant, Trey 0.0
LR20200824 - Zhang, Yang 40.0
LR20200824 - Zhang, Xianyu 54.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

DE

  • Startup!
  • Repository NFS export fix
  • Newdata fix
    • Ultimately swapped VMDK for a much larger one
    • Cristina still working on smoothing out all aspects of data flow (ACL issue)
  • Summit mount fixes after swapping newdata disk
  • TFTP server config on archive-in
  • XZ weird postfix problem on VM, fixed by changing localhost to 127.0.0.1 in postfix config
  • GR Adobe install
  • Jim Solidworks local install--wip
  • INAF VPN account--make test accounts in Affiliates group to see what's up with DNS--wip
  • Resolved 4D machine ID blackbox--no idea what they use to generate it but moving the VM was enough to change it
  • tcs machines NFS mount only responding to * export--wip
    • Open ticket with Synology to see if they can find evidence in the logs that are hidden from me
  • web.tucson not mounting some homes fixed but needs to be discussed with SH re: systemd mount vs. fstab
  • Chasing Seagate RMA, one 8TB drive on its way to me.
  • Opened conversation with Carlos re: possible server room power work
Next:
  • Reconfigure Vault backup disk to occupy less space
  • as-you-can work list from IT meeting notes
  • Fires, hopefully only figurative
  • Syno support wants me to log the broken mounts. They asked on Friday evening during night ops, so I'll use the test mount I made.

PK

YZ

  • Previous:
    • Some LBC work:
      • Reviewed some LBC code following the /newdata problem.
      • Preparing for new bias level update (hasn't happened yet due to problems with LBCB)
    • Some TCS-related exploration, mainly for PMC and slew/preset simulations. John Hill provided some helpful tips.
    • TMS work (meeting, wiki work, understanding our new design with regards to missing channel information).
    • AO-Cravat work (meeting, wiki work, exploring FWHM and releated computations).
  • Next:
    • Possible LBC work.
    • Also Olga brought up the topic of proper/better LBC conf management (worth a discussion).
    • General TMS and AO-Cravat work (lots of things on the plate here).

CH

  • Last week:
    • AGW1 testing support.
    • oacontrol: Fixed typo preventing compilation on Centos7/64 bit. Addressed new typo per PK comment on my 5.2.1 patch.
    • LBTI alert script: fixed source of byte vs. str error in pyindi.py, getting good data.
    • Develop oacontrol error improvements.
    • Replaced workstation power supply.
  • This week:
    • LBTI alert script: finish release of script.
    • Fix oacontrol PR.

TB

  • Last Week:
    • Program now prints start date with every next/prev command
    • Some progress with resolving wind direction 0=360 issue, using np.unwrap(direction_data_in_radians)
      • Might need to look at a different avenue/more manual solution
    • Some more reading
    • COVID Training
  • Next Week:
    • More reading
    • Better solutions for wind direction
    • More next steps

MB

  • Last week
    • OC, Management
    • mech interview #2
    • PE's one on one and completion
    • ATT draft KoM and collection documentation
    • Wired registration follow-up (...today)
    • Misc IT and archive support
  • This week
    • PE's very last batch
    • ATT
    • AAS job regiser ad
    • DMS websockets, AO, AdSec, etc. integrate demo ui, LBC stiched png preview, max_bytes for ovms

XZ

  • Last week
    • DX AdSec issue document to Christian
    • SX AdSec IT#8168 analysis and document
    • AO support, DSP watchdog expiration discussion, contact 4D etc
    • Meetings
  • This week
    • Contiue AdSec related
    • script for moving ASM and WFS log and telemetry files
    • Safety training
    • AO support

IS

Last week:
  • Support
  • PIT
    • Most of the week dedicated to this work. I had to find new logic in order to include PEPSI as a pairing instrument w/ binocular observations. Still working on it.
  • LUCI
    • Mask upload and CCR WEBIO signal reading
This week:
  • Safety Training
  • LUCI
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
    • Finish document and configuration for WebIO
  • PIT
    • Finish PIT release.
    • Work on ways to apply an RPM to it, even though the PIT might not be the best example application but it should be very similar to
the OT.
Topic revision: r15 - 31 Aug 2020, CarlHolmberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback